Nadine is two years younger than Mark. The sum of their ages is 28. How old is Nadine?
Answer by ikleyn(52847)

.

Nadine is 2 years younger than Mark  <------>  Mark is 2 year older than Nadine.



Nadine + Mark        = 28


Nadine + (Nadine +2) = 28


2*Nadine = 28-2 = 26


Nadine = 26/2 = 13 years old.   ANSWER